OpenStudy (freemap):

Last week, Hazel paid $32 to have 50 reports printed. This week, she paid $57 to have 100 reports printed. Assume the relationship is linear. Which equation models the relationship between the number of reports, x, and the cost, y? A. y=2x+68 B. y=1/2x +7 C. y=1/2x+24 D. y=2x +34

OpenStudy (freemap):

How do I solve this?

Vocaloid (vocaloid):

x = number of reports y = cost using the information in the problem, we can write two points (32,50) and (57,100) then you can write an equation for the line that goes through these points

OpenStudy (freemap):

ok, can you stay while I solve it

Vocaloid (vocaloid):

sure

Vocaloid (vocaloid):

one more time: (50,32) and (100,57)

OpenStudy (freemap):

lol ok

OpenStudy (freemap):

If you use slope, its .418

Vocaloid (vocaloid):

you sure? (50,32) and (100,57) slope = (y2-y1)/(x2-x1)

Vocaloid (vocaloid):

(57-32)/(100-25) = ?

OpenStudy (freemap):

ok wow I did that wrong, sorry

OpenStudy (freemap):

-7/43= -.1627

Vocaloid (vocaloid):

still not right

Vocaloid (vocaloid):

57-32 = ?

OpenStudy (freemap):

25

Vocaloid (vocaloid):

100-50 = ?

OpenStudy (freemap):

oh, i thought i had to use slope

Vocaloid (vocaloid):

we are using slope

Vocaloid (vocaloid):

(57-32)/(100-50) = ?

OpenStudy (freemap):

25/50 =.5=1/2

Vocaloid (vocaloid):

good, so our slope is 1/2 know how to finish from here?

OpenStudy (freemap):

I have no clue how to finish

OpenStudy (freemap):

please help me out

OpenStudy (freemap):

B or C

Vocaloid (vocaloid):

use the point-slope equation y-y1 = m(x-x1) where (x1,y1) = one of the points (either one, take your pick) and m = slope

OpenStudy (freemap):

OK, just a sec

OpenStudy (freemap):

y-32=1/2(x-50)

OpenStudy (freemap):

32+-50 is 18

OpenStudy (freemap):

oh wait do I have to multiply 1/2 times -50

OpenStudy (freemap):

Please help

OpenStudy (freemap):

@ganeshie8

OpenStudy (freemap):

B or C

Vocaloid (vocaloid):

y-32=1/2(x-50) start by distributing the 1/2

OpenStudy (freemap):

I got 7 so the answer is B, forgot to distribute. Thanks so much, real hero